# Taking quantisation seriously: a farewell to waves

### v2

@@ 8 Conclusion
The de Broglie wavelength is not covariant as the length scale of a simultaneous extended structure in any reference frame. Thus, we cannot use it as physical explanation of observed phenomena, it is merely a useful heuristic. This puts claims about wave-particle duality under extreme strain and undermines otherwise compelling explanations of interference and diffraction. To resolve this, we have argued that the apparent wave nature of quantum objects is only manifest when diffraction elements are treated classically. In contrast to the Kapitza-Dirac effect, which never needs to invoke either the waviness of incident electrons or the scattering photons. When the diffraction element is considered as a participating quantum system, a picture of diffraction and interference emerges that describes them as apparent consequences of quantised momentum exchanges. Thus, uniting phase-grating and barrier diffraction experiments with one physically consistent explanation. This is justified by the periodic structure of the scatterer implying it must have quantised momentum eigenstates. Moreover, a double slit analogue with the explicitly quantised photon orbital angular momentum shows the same expected distribution as the linear double slit momentum. This provides evidence that a quantised exchange interaction can produce what looks like an interference pattern. Moreover, the only conclusion to be drawn from the analogy is that the interaction is responsible for quantising linear momentum in a double slit. Vitally, it is shown that the momentum distribution is determined at the slits themselves, not at the point of particle detection. Finally, we also presented evidence that this viewpoint is already used implicitly by experimenters making use of electron diffraction to characterise phonons in condensed matter and this is the only physically consistent way to view these experiments.
This work demonstrates that complex amplitude interference is capable of representing particle physics that it does not resemble at all. Thus, explaining why naive positivism resulted in a view of quantum mechanics at odds with realism and logic. Moreover, we see an indication of a relationship between quantisation and apparent wave behaviour being mediated via complex amplitude interference, meaning that the heuristic value of waviness is not just a deceptive coincidence.
It is not immediately clear how this viewpoint generalises to interference phenomena related to beam splitters. Such a study would hinge upon a description of beam-splitters as active participants in interactions with photons, something beyond the scope of this work. However, our conclusions here raise the possibility that a similar realist interpretation could be buried beneath the obscuring mathematics.
@@ Acknowledgements
Thanks to Robert de Mello Koch, Sam van Leuven, and Andrew Forbes for helpful discussions; as well as Emil Roduner and Tjaart Kruger for their encouragement.

### v4

@@ 8 Conclusion
The de Broglie wavelength is not covariant as the length scale of a simultaneous extended structure in any reference frame. Thus, we cannot use it as physical explanation of observed phenomena, it is merely a useful heuristic. This puts claims about wave-particle duality under extreme strain and undermines otherwise compelling explanations of interference and diffraction.
@@
To resolve this, we have argued that the apparent wave nature of quantum objects is only manifest when diffraction elements are treated classically [14]. In contrast to the Kapitza-Dirac effect, which never needs to invoke either the waviness of incident electrons or the scattering photons. When the diffraction element is considered as a participating quantum system, a picture of diffraction and interference emerges that describes them as apparent consequences of quantised momentum exchanges. Thus, uniting phase-grating and barrier diffraction experiments with one physically consistent explanation. This is justified by the periodic structure of the scatterer implying it must have quantised momentum eigenstates.
We have also presented experiments that signal the validity of our explanation of waviness: ultrafast electron diffraction, where phonon excitation alters diffraction patterns, and polarisable molecule diffraction, with a degeneracy between slit width and interaction strength. Further points of interest were the lack of a coherence requirement in diffraction experiments, in favour of collimation instead [44], and indications of momentum conservation on a per event level [43, 49, 51]. Moreover, a double slit analogue with the explicitly quantised photon orbital angular momentum shows the same expected distribution as the linear double slit momentum. This provides evidence that a quantised exchange interaction can produce what looks like an interference pattern. Further, the only conclusion to be drawn from the analogy is that the interaction is responsible for quantising linear momentum in a regular double slit.
We revisited the double slit and demonstrated the momentum distribution is determined at the slits themselves, not at the point of particle detection and before wave overlap could even occur. In addition, it does not change during propagation to the screen. Combined with momentum conservation [43] and issues of coherence [44], this is a powerful argument in favour of discarding wave overlap as a physical explanation. Finally, we broadened this discussion to include “which-way” information. Showing that this can even be shorn of mystical associations in the process. In particular, we analysed a “quantum eraser” to demonstrate that interference patterns always exist in this scenario, they simply occur in complementary pairs as a consequence of parity symmetry being broken by polarisers.
This work demonstrates that complex amplitude interference is capable of representing particle physics that it does not resemble at all. Thus, explaining why naive positivism resulted in a view of quantum mechanics at odds with realism and logic. Interestingly, the source of both the quantisation and apparent waviness is symmetry, explaining why we can employ an inconsistent physical picture but obtain the correct predictions anyway. This link also explains why group theory has such utility in quantum mechanics: the tools of symmetry are naturally useful to describe quantised systems where the quantisation itself arises from symmetries.
While we do not cover interference with beam splitters, arguments have already been provided that these cases must be understood in terms of interactions between particle and beam splitter that involve exchange of energy and momentum [14]. This is entirely in keeping with the main thesis here, that we should not regard apparatus as inert in quantum experiments.
@@ Acknowledgements
Thanks to Andrew Forbes, Thomas Konrad, Robert de Mello Koch, and Sam van Leuven for helpful discussions; as well as Emil Roduner and Tjaart Kruger for their encouragement.